OpenStudy (anonymous):

Solve y = 3x – 12 by the substitution method. 2x – 3y = -13

OpenStudy (anonymous):

2x - 3y = -13. Substitute the value of y = 3x - 12 from the equation1. 2x -3*(3x - 12) = -13 2x - 9x + 36 = -13 x =...